if csc(theta) =2.5 find sec(pi/2- theta)? Would you please describe your steps?
Answer by Alan3354(31538)

if csc(theta) =2.5 find sec(pi/2- theta)? Would you please describe your steps?
--------------------
sin(x) = cos(pi/2-x) --> csc(x) = sec(pi/2-x) and it's done.
sec(pi/2 - theta) = 2.5